# The Uniform Mordell–Lang Conjecture

### v3

@@ 5.1. A result in family
Let $\bar{\eta}$ be the geometric generic point of $S^{\prime}$. Invoke [Zha98, proof of Lem.3.1] to $\mathscr{X}_{\bar{\eta}}^{m}\subseteq\mathcal{A}_{\bar{\eta}}^{m}$. Then there exists an $M_{0}(S^{\prime})>0$ with the following property: For each $M\geq M_{0}(S^{\prime})$, a generic fiber of $\mathscr{D}_{M}^{\mathcal{A}_{S^{\prime}}^{[m]}}|_{\mathscr{X}_{\bar{\eta}}^{m(M+1)}}$ is a $\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}}(\mathscr{X}_{\bar{\eta}})^{m}$-orbit; here $\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}}(\mathscr{X}_{\bar{\eta}})^{m}=\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}^{m}}(\mathscr{X}_{\bar{\eta}}^{m})$ is viewed as a subgroup of $(\mathcal{A}_{\bar{\eta}}^{m})^{M+1}$ via the diagonal embedding. Notice that $\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}}(\mathscr{X}_{\bar{\eta}})^{m}$ is finite by definition of $S_{\mathrm{gen}}$.
Take $M=\max_{S^{\prime}}\{M_{0}(S^{\prime})\}>0$ with $S^{\prime}$ running over all irreducible components of $S_{\mathrm{gen}}$. Then $M=M(S)>0$ depends only on $S$.
We has an $S^{\prime}$-morphism $(\mathrm{id},\mathscr{D}_{M}^{\mathcal{A}_{S^{\prime}}^{[m]}})\colon\mathcal{A}_{S^{\prime}}^{[m]}\times_{S^{\prime}}(\mathcal{A}_{S^{\prime}}^{[m]})^{[M+1]}\rightarrow\mathcal{A}_{S^{\prime}}^{[m]}\times_{S^{\prime}}(\mathcal{A}_{S^{\prime}}^{[m]})^{[M]}$. Consider its restriction to $\mathscr{X}^{[m]}_{S^{\prime}}\times_{S^{\prime}}(\mathscr{X}^{[m]}_{S^{\prime}})^{[M+1]}=\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$
(5.6) $\mathscr{D}:=(\mathrm{id},\mathscr{D}_{M}^{\mathcal{A}_{S^{\prime}}^{[m]}})|_{\mathscr{X}^{[m(M+2)]}_{S^{\prime}}}\colon\mathscr{X}^{[m(M+2)]}_{S^{\prime}}=\mathscr{X}^{[m]}_{S^{\prime}}\times_{S^{\prime}}(\mathscr{X}^{[m]}_{S^{\prime}})^{[M+1]}\rightarrow\mathcal{A}_{S^{\prime}}^{[m]}\times_{S^{\prime}}(\mathcal{A}_{S^{\prime}}^{[m]})^{[M]}=\mathcal{A}_{S^{\prime}}^{[m(M+1)]}.$
In $\mathcal{A}_{S^{\prime}}^{[m(M+2)]}$, we have a non-degenerate subvariety $\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$. Moreover, $\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})$ is non-degenerate in $\mathcal{A}_{S^{\prime}}^{[m(M+1)]}=\mathcal{A}_{S^{\prime}}^{[m]}\times_{S^{\prime}}(\mathcal{A}_{S^{\prime}}^{[m]})^{[M]}$. Indeed,
$\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}))=\mathscr{D}(\mathscr{X}^{[m]}_{S^{\prime}}\times_{S^{\prime}}(\mathscr{X}^{[m]}_{S^{\prime}})^{[M+1]})=\mathscr{X}^{[m]}_{S^{\prime}}\times_{S^{\prime}}\mathscr{D}_{M}^{\mathcal{A}_{S^{\prime}}^{[m]}}((\mathscr{X}^{[m]}_{S^{\prime}})^{[M+1]})\subseteq\mathcal{A}_{S^{\prime}}^{[m]}\times_{S^{\prime}}(\mathcal{A}_{S^{\prime}}^{[m]})^{[M]},$
and hence is non-degenerate because $\mathscr{X}^{[m]}_{S^{\prime}}$ is non-degenerate; see [Gao21, Lem.6.2].
Now we have obtained the two desired non-degenerate subvarieties $\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$ in $\mathcal{A}_{S^{\prime}}^{[m(M+2)]}$ and $\mathscr{D}(\mathscr{X}^{[m(M+2]}_{S^{\prime}})$ in $\mathcal{A}_{S^{\prime}}^{[m(M+1)]}$. Moreover using the notations below (5.5), a generic fiber of $\mathscr{D}|_{\mathscr{X}_{\bar{\eta}}^{m(M+2)}}$ is a $G_{\bar{\eta}}:=\{0\}\times\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}}(\mathscr{X}_{\bar{\eta}})^{m}$-orbit. Moreover, $G_{\bar{\eta}}$ is a finite group and hence extends to a finite étale $S^{\prime}$-group subscheme $G$ of $\mathcal{A}_{S^{\prime}}^{[m(M+2)]}\rightarrow S^{\prime}$.
Step 2 Choose suitable functions $f_{S^{\prime},1},f_{S^{\prime},2}$ and constant $\epsilon_{S^{\prime}}>0$ for later applications of equidistribution.
Write $\omega_{m(M+2)}$ for the Betti form on $\mathcal{A}_{S^{\prime}}^{[m(M+2)],\mathrm{an}}$ and $\omega_{m(M+1)}$ for the Betti form on $\mathcal{A}_{S^{\prime}}^{[m(M+1)],\mathrm{an}}$.
Let $\mu_{S^{\prime},1}$ be the measure on $\mathscr{X}^{[m(M+2)]}_{S^{\prime}}(\mathbb{C})$ as in [Küh21, Thm.1]; it equals $k_{S^{\prime},1}\cdot\omega_{m(M+2)}|_{\mathscr{X}^{[m(M+2)]}_{S^{\prime}}(\mathbb{C})}^{\wedge\dim\mathscr{X}^{[m(M+2)]}_{S^{\prime}}}$ with a constant $k_{S^{\prime},1}>0$.
Let $\mu_{S^{\prime},2}$ be the measure on $\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})(\mathbb{C})$ as in [Küh21, Thm.1]; it equals $k_{S^{\prime},2}\cdot\omega_{m(M+1)}|_{\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})(\mathbb{C})}^{\wedge\dim\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})}$ with a constant $k_{S^{\prime},2}>0$.
We start by proving $\mu_{S^{\prime},1}\not=\mathscr{D}^{*}\mu_{S^{\prime},2}$.
For the point $\mathbf{x}\in(\mathscr{X}^{[m]}_{S^{\prime}})^{\mathrm{sm}}(\mathbb{C})$ from (5.4), denote by $\Delta_{\mathbf{x}}$ the point $(\mathbf{x},\ldots,\mathbf{x})$ in $(\mathscr{X}^{[m]}_{S^{\prime}})^{[M+1]}(\mathbb{C})=\mathscr{X}^{[m(M+2)]}_{S^{\prime}}(\mathbb{C})$. Then $(\mathbf{x},\Delta_{\mathbf{x}})\in(\mathscr{X}^{[m]}_{S^{\prime}}\times_{S^{\prime}}(\mathscr{X}^{[m]}_{S^{\prime}})^{[M+1]})(\mathbb{C})$, which is furthermore a smooth point our choice of $\mathbf{x}$ below (5.4). We have $(\mu_{S^{\prime},1})_{(\mathbf{x},\Delta_{\mathbf{x}})}\not=0$; see [Gao21, Lem.6.3].
On the other hand, $\mathscr{D}_{M}^{\mathcal{A}_{S^{\prime}}^{[m]}}(\Delta_{\mathbf{x}})$ is the origin of fiber of $(\mathcal{A}_{S^{\prime}}^{[m]})^{[M]}\rightarrow S^{\prime}$ in question (which we call $(\mathcal{A}_{s}^{m})^{M}$), so $\mathscr{D}_{M}^{\mathcal{A}_{S^{\prime}}^{[m]}}|_{\mathscr{X}^{[m(M+1)]}_{S^{\prime}}}^{-1}(\mathscr{D}_{M}^{\mathcal{A}_{S^{\prime}}^{[m]}}(\Delta_{\mathbf{x}}))$ contains the diagonal of $\mathscr{X}_{s}^{m}\subseteq\mathcal{A}_{s}^{m}$ in $(\mathcal{A}_{s}^{m})^{M}$ (which for the moment we denote by $\Delta_{\mathscr{X}_{s}^{m}}$). Therefore for the morphism $\mathscr{D}=(\mathrm{id},\mathscr{D}_{M}^{\mathcal{A}_{S^{\prime}}^{[m]}})|_{\mathscr{X}^{[m(M+2)]}_{S^{\prime}}}$ from (5.6), $\mathscr{D}^{-1}(\mathscr{D}(\mathbf{x},\Delta_{\mathbf{x}}))$ contains $(\mathbf{x},\Delta_{\mathscr{X}_{s}^{m}})$. Thus $\dim\mathscr{D}^{-1}(\mathscr{D}(\mathbf{x},\Delta_{\mathbf{x}}))>0$, so the linear map
$\mathrm{d}\mathscr{D}\colon T_{(\mathbf{x},\Delta_{\mathbf{x}})}\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\rightarrow T_{\mathscr{D}(\mathbf{x},\Delta_{\mathbf{x}})}\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})$
has non-trivial kernel. This implies $(\mathscr{D}^{*}\mu_{S^{\prime},2})_{(\mathbf{x},\Delta_{\mathbf{x}})}=0$.
Thus we get $\mu_{S^{\prime},1}\not=\mathscr{D}^{*}\mu_{S^{\prime},2}$ by looking at their evaluations at $(\mathbf{x},\Delta_{\mathbf{x}})$.
Thus there exist a constant $\epsilon_{S^{\prime}}>0$ and a function $f_{S^{\prime},1}\in\mathscr{C}_{\mathrm{c}}^{0}(\mathscr{X}^{[m(M+2)],\mathrm{an}}_{S^{\prime}})$ such that
(5.7) $\left|\int_{\mathscr{X}^{[m(M+2)],\mathrm{an}}_{S^{\prime}}}f_{S^{\prime},1}\mu_{S^{\prime},1}-\int_{\mathscr{X}^{[m(M+2)],\mathrm{an}}_{S^{\prime}}}f_{S^{\prime},1}\mathscr{D}^{*}\mu_{S^{\prime},2}\right|>\epsilon_{S^{\prime}}.$
We finish this step by showing that we can choose such an $f_{S^{\prime},1}$ with the following property: there exists a unique $f_{S^{\prime},2}\in\mathscr{C}_{\mathrm{c}}^{0}(\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})^{\mathrm{an}})$ such that $f_{S^{\prime},1}=f_{S^{\prime},2}\circ\mathscr{D}$.
@@
Indeed, let $G$ be the finite $S^{\prime}$-group from the end of Step 1. Then our $\mathscr{D}$ from (5.6) satisfies the following property. There exists a Zariski open dense subset $V$ of $\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})\subseteq\mathcal{A}_{S^{\prime}}^{[m(M+1)]}$ such that $\mathscr{D}|_{\mathscr{D}^{-1}(V)}:\mathscr{D}^{-1}(V)\rightarrow V$ is finite étale and that each of its fibers $\mathscr{D}^{-1}(y)$ is a $G_{s}$-orbit for each $y\in V(\mathbb{C})$; here $y\mapsto s\in S^{\prime}(\mathbb{C})$ under $\mathcal{A}_{S^{\prime}}^{[m(M+1)]}\rightarrow S^{\prime}$ and $G_{s}$ means the fiber of $G\rightarrow S^{\prime}$ over $s$. Moreover, we can assume that $V$ is sufficiently small such that each point $\sigma\in G_{\bar{\eta}}$ extends to a section $\sigma_{U}\colon U\rightarrow G|_{U}\subset\mathcal{A}_{U}^{[m(M+2)]}$ where $U\subseteq S^{\prime}$ is the image of $V$. We may and do assume that $f_{S^{\prime},1}$ is supported in $\mathscr{D}^{-1}(V)$. For each $\sigma\in G_{\bar{\eta}}$, by abuse of notation we also use $\sigma_{U}$ to denote the translation of $\mathcal{A}_{U}^{[m(M+2)]}$ defined by it. As $\sigma$ is the $(M+1)$-fold self-product of a translation of $\mathcal{A}_{S^{\prime}}^{[m]}$, we have $\mathscr{D}\circ\sigma=\mathscr{D}$ and $f_{S^{\prime},1}^{*}:=\sum_{\sigma\in G_{\overline{\eta}}}f_{S^{\prime},1}\circ\sigma_{U}$ equals $f_{S^{\prime},2}\circ\mathscr{D}$ for an $f_{S^{\prime},2}\in\mathscr{C}_{\mathrm{c}}^{0}(\mathscr{D}(V))$. Notice that $\sigma_{U}^{*}\mu_{S^{\prime},1}=\mu_{S^{\prime},1}$ over $U$ for each $\sigma\in G_{\bar{\eta}}$ because $\mu_{S^{\prime},1}$ is obtained from the Betti form. Thus (5.7) holds true if we replace $f_{S^{\prime},1}$ by $f_{S^{\prime},1}^{*}$, and so it suffices to replace $f_{S^{\prime},1}$ by $f_{S^{\prime},1}^{*}$.
Step 3 Prove some height lower bound on $\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$ or $\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})$.
We apply the third-named author’s equidistribution result [Küh21, Thm.1] twice. In fact, to prove the desired lower bound it is more convenient to apply its corollary under the formulation of [Gao21, Cor.8.2].
@@
Apply [Gao21, Cor.8.2] to $\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$, $f_{S^{\prime},1}$ and $\epsilon_{S^{\prime}}$. We thus obtain a constant $\delta_{\epsilon_{S^{\prime}},1}>0$ and a Zariski closed proper subset $Z_{S^{\prime},1}:=Z_{f_{S^{\prime},1},\epsilon_{S^{\prime}}}$ of $\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$. Apply [Gao21, Cor.8.2] to $\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})$, $f_{S^{\prime},2}$ and $\epsilon_{S^{\prime}}$. We thus obtain a constant $\delta_{\epsilon_{S^{\prime}},2}>0$ and a Zariski closed proper subset $Z_{S^{\prime},2}:=Z_{f_{S^{\prime},2},\epsilon_{S^{\prime}}}$ of $\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})$.
Let $\delta_{S^{\prime}}:=\min\{\delta_{\epsilon_{S^{\prime}},1},\delta_{\epsilon_{S^{\prime}},2}\}>0$, and let $Z_{S^{\prime}}=Z_{S^{\prime},1}\bigcup\mathscr{D}^{-1}(Z_{S^{\prime},2})\bigcup Z_{S^{\prime},3}$, where $Z_{S^{\prime},3}$ is the largest Zariski closed subset of $\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$ on which $\mathscr{D}$ is not finite. Then $Z_{S^{\prime}}$ is Zariski closed in $\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$, and is proper because $\mathscr{D}$ is generically finite. If a point $\mathbf{x}\in(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\setminus Z_{S^{\prime}})(\overline{\mathbb{Q}})$ is such that $\hat{h}_{{\mathcal{L}}^{\boxtimes m(M+2)}}(\mathbf{x})<\delta_{S^{\prime}}$ and $\hat{h}_{{\mathcal{L}}^{\boxtimes m(M+1)}}(\mathscr{D}(\mathbf{x}))<\delta_{S^{\prime}}$, then case (ii) of [Gao21, Cor.8.2] holds true for both $\mathbf{x},f_{S^{\prime},1},\mu_{S^{\prime},1}$ and $\mathscr{D}(\mathbf{x}),f_{S^{\prime},2},\mu_{S^{\prime},2}$. Thus
@@
$\left|\int_{\mathscr{X}^{[m(M+2)],\mathrm{an}}_{S^{\prime}}}f_{S^{\prime},1}\mu_{S^{\prime},1}-\frac{1}{\#O(\mathbf{x})}\sum_{y\in O(\mathbf{x})}f_{S^{\prime},1}(y)\right|<\epsilon_{S^{\prime}}\text{ and }\left|\int_{\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})^{\mathrm{an}}}f_{S^{\prime},2}\mu_{S^{\prime},2}-\frac{1}{\#O(\mathscr{D}(\mathbf{x}))}\sum_{y\in O(\mathscr{D}(\mathbf{x}))}f_{S^{\prime},2}(y)\right|<\epsilon_{S^{\prime}}$
where $O(\cdot)$ is the Galois orbit. But $\frac{1}{\#O(\mathbf{x})}\sum_{y\in O(\mathbf{x})}f_{S^{\prime},1}(y)=\frac{1}{\#O(\mathscr{D}(\mathbf{x}))}\sum_{y\in O(\mathscr{D}(\mathbf{x}))}f_{S^{\prime},2}(y)$ because $f_{S^{\prime},1}=f_{S^{\prime},2}\circ\mathscr{D}$. So we have
@@
$\left|\int_{\mathscr{X}^{[m(M+2)],\mathrm{an}}_{S^{\prime}}}f_{S^{\prime},1}\mu_{S^{\prime},1}-\int_{\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})^{\mathrm{an}}}f_{S^{\prime},2}\mu_{S^{\prime},2}\right|\leq 2\epsilon_{S^{\prime}}.$
This contradicts (5.7) because $f_{S^{\prime},1}=f_{S^{\prime},2}\circ\mathscr{D}$.
@@
Hence for each point $\mathbf{x}\in(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\setminus Z_{S^{\prime}})(\overline{\mathbb{Q}})$, we are in one of the following alternatives.

### v4

@@ 5.1. A preliminary result for the restricted Hilbert scheme
Let $\bar{\eta}$ be the geometric generic point of $S^{\prime}$. By the proof of [Zha98, Lem. 3.1], there exists $M_{0}(S^{\prime})>0$ with the following property: For each $M\geq M_{0}(S^{\prime})$, the generic fiber of the restriction of $\mathscr{D}_{0}$ to $\mathscr{X}_{\bar{\eta}}^{m(M+1)}$ is an orbit under the action of $\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}^{m}}(\mathscr{X}_{\bar{\eta}}^{m})\subseteq\mathcal{A}_{\bar{\eta}}^{m}$ diagonally embedded into $(\mathcal{A}_{\bar{\eta}}^{m})^{M+1}$. Notice that $\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}^{m}}(\mathscr{X}_{\bar{\eta}}^{m})=\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}}(\mathscr{X}_{\bar{\eta}})^{m}$ is finite by the definition of $S_{\mathrm{gen}}$. We set $M=\max_{S^{\prime}}\{M_{0}(S^{\prime})\}>0$ where $S^{\prime}$ runs over all irreducible components of $S_{\mathrm{gen}}$. Then $M=M(S)>0$ depends only on $S$.
We further define the $S^{\prime}$-morphism
$\mathscr{D}_{0}^{\prime}:=(\mathrm{id},\mathscr{D}_{0})\colon\ \ \mathcal{A}_{S^{\prime}}^{[m]}\times_{S^{\prime}}(\mathcal{A}_{S^{\prime}}^{[m]})^{[M+1]}\longrightarrow\mathcal{A}_{S^{\prime}}^{[m]}\times_{S^{\prime}}(\mathcal{A}_{S^{\prime}}^{[m]})^{[M]}$
and its restriction
(5.6) $\mathscr{D}:\ \ \mathscr{X}^{[m]}_{S^{\prime}}\times_{S^{\prime}}(\mathscr{X}^{[m]}_{S^{\prime}})^{[M+1]}=\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\longrightarrow\mathscr{D}_{0}^{\prime}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})$.
(Note that we also (co-)restrict on the range to simplify our notation.) For later use, we set
$G_{\bar{\eta}}:=\{0\}\times\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}}(\mathscr{X}_{\bar{\eta}})^{m}\subseteq\mathcal{A}_{\bar{\eta}}^{m}\times(\mathcal{A}_{\bar{\eta}}^{m})^{M+1}=\mathcal{A}_{\bar{\eta}}^{m(M+2)};$
then the fibers of $\mathscr{D}|_{\overline{\eta}}$ are naturally orbits under $G_{\overline{\eta}}$.
The subvarieties
$\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\subseteq\mathcal{A}_{S^{\prime}}^{[m(M+2)]}\ \ \text{and}\ \ \mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})\subseteq\mathcal{A}_{S^{\prime}}^{[m(M+1)]}$
are both non-degenerate. For the former one, this follows directly from our choice of $m$. For the latter one, we remark that
$\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}))=\mathscr{X}^{[m]}_{S^{\prime}}\times_{S^{\prime}}\mathscr{D}_{0}((\mathscr{X}^{[m]}_{S^{\prime}})^{[M+1]}),$
whose non-degeneracy follows from the non-degeneracy of $\mathscr{X}^{[m]}_{S^{\prime}}$; see for example [Küh21, Lem. 24] or [Gao21, Lem. 6.2].
Step 2: Defining non-proportional measures $\mu_{S^{\prime},1}$ and $\mathscr{D}^{*}\mu_{S^{\prime},2}$ on $\mathcal{A}_{S^{\prime}}^{[m(M+2)]}(\mathbb{C})$
[…]
$(\omega_{m(M+2)})^{\wedge(rm(M+2)+\dim(S^{\prime}))}\ \ \text{(resp.\ }(\omega_{m(M+1)})^{\wedge(rm(M+2)+\dim(S^{\prime}))}).)$
We use this to prove that the measures $\mu_{S^{\prime},1}$ and $\mathscr{D}^{*}\mu_{S^{\prime},2}$ are non-proportional. In particular, $\mu_{S^{\prime},1}\not=\mathscr{D}^{*}\mu_{S^{\prime},2}$.
[…]
for its $(M+1)$-fold self product. For the point $\mathbf{z}\in(\mathscr{X}^{[m]}_{S^{\prime}})^{\mathrm{sm}}(\mathbb{C})$ chosen above (5.4), the point $(\mathbf{z},\Delta_{\mathbf{z}})$ is a smooth point of $(\mathscr{X}^{[m]}_{S^{\prime}}\times_{S^{\prime}}(\mathscr{X}^{[m]}_{S^{\prime}})^{[M+1]})(\mathbb{C})$ and
$(\mu_{S^{\prime},1})_{(\mathbf{z},\Delta_{\mathbf{z}})}\not=0;$
see for example [Küh21, Lem. 11 and 25].
By definition, we have $\mathscr{D}((\mathbf{z},\Delta_{\mathbf{z}}))=(\mathbf{z},0,\dots,0)$ and hence
[…]
As this is of dimension $rm+\dim S^{\prime}>0$ locally at $(\mathbf{z},\Delta_{\mathbf{z}})$, the differential
$d\mathscr{D}\colon T_{(\mathbf{z},\Delta_{\mathbf{z}})}\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\longrightarrow T_{\mathscr{D}(\mathbf{z},\Delta_{\mathbf{z}})}\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})$
has a non-trivial kernel. So
$(\mathscr{D}^{*}\mu_{S^{\prime},2})_{(\mathbf{z},\Delta_{\mathbf{z}})}=0,$
and, consequently, $\mu_{S^{\prime},1}\not=\mathscr{D}^{*}\mu_{S^{\prime},2}$.
Step 3: Choice of test functions $f_{S^{\prime},1},f_{S^{\prime},2}$ and $\varepsilon_{S^{\prime}}>0$
[…]
(5.7) $\left|\int_{\mathscr{X}^{[m(M+2)]}_{S^{\prime}}(\mathbb{C})}f_{S^{\prime},1}\mu_{S^{\prime},1}-\int_{\mathscr{X}^{[m(M+2)]}_{S^{\prime}}(\mathbb{C})}f_{S^{\prime},1}\mathscr{D}^{*}\mu_{S^{\prime},2}\right|>2\varepsilon_{S^{\prime}}.$
This is, however, not enough for our purposes. We want $f_{S^{\prime},1}$ to be of the form
$f_{S^{\prime},1}=f_{S^{\prime},2}\circ\mathscr{D}$
[…]
@@
To achieve this goal, let $G_{\overline{\eta}}$ be the finite group defined in Step 1. By abuse of notation denote the projection by $\pi\colon\mathcal{A}_{S^{\prime}}^{[m(M+2)]}\rightarrow S^{\prime}$. There exists a Zariski open dense $U\subseteq S^{\prime}$ such that $G_{\overline{\eta}}$ extends to a flat group scheme $G$ over $U$. Furthermore, there exists a Zariski open dense subset $V\subseteq\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{U})$ such that $\mathscr{D}|_{\mathscr{D}^{-1}(V)}\colon\mathscr{D}^{-1}(V)\rightarrow V$ is finite étale and each of its fibers $\mathscr{D}^{-1}(\mathbf{y})$, $\mathbf{y}\in V(\mathbb{C})$, is a $G_{\pi(\mathbf{y})}$-orbit, where $G_{\pi(\mathbf{y})}$ is the restriction of $G$ to the fiber over $\pi(\mathbf{y})$. We may and do assume that $f_{S^{\prime},1}$ is supported in $V\subseteq\pi^{-1}(U)$ without compromising (5.7). Furthermore, we can replace $f_{S^{\prime},1}$ by
$f_{S^{\prime},1}\colon\ \$$\mathscr{X}^{[m(M+2)]}_{U} \displaystyle\longrightarrow \displaystyle\mathbb{R}$
$(\mathbf{z}_{0},\mathbf{z}_{1},\ldots,\mathbf{z}_{M})$ $\longmapsto$ $\sum_{\mathbf{t}\in G_{s}}f_{S^{\prime},1}(\mathbf{z}_{0}+\mathbf{t},\dots,\mathbf{z}_{M}+\mathbf{t})$
where $G_{s}$ is the restriction of $G$ to the fiber over $s=\pi(\mathbf{z}_{0},\dots,\mathbf{z}_{M})$ and (5.7) remains valid; indeed, both $\mu_{S^{\prime},1}$ and $\mu_{S^{\prime},2}$ are translation invariant since they are exterior powers of Betti forms, which are translation invariant.
In summary, $f_{S^{\prime},1}$ is constant on the fibers of $\mathscr{D}$. Since its support is contained in the locus where $\mathscr{D}$ is étale, there exists a unique function $f_{S^{\prime},2}$ such that $f_{S^{\prime},1}=f_{S^{\prime},2}\circ\mathscr{D}$.
Step 4: Obtaining a height lower bound via equidistribution
@@
We apply the equidistribution result [Küh21, Thm. 1] in the form of [Küh21, Lem. 23] twice. Applied to $\mathscr{X}_{S^{\prime}}^{[m(M+2)]}$, $f_{S^{\prime},1}$ and $\varepsilon_{S^{\prime}}$, we obtain a constant $\delta_{\varepsilon_{S^{\prime}},1}>0$ and a Zariski closed subset $Z_{S^{\prime},1}\subsetneq\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$ such that
@@
(5.8) $\left|\int_{\mathscr{X}^{[m(M+2)]}_{S^{\prime}}(\mathbb{C})}f_{S^{\prime},1}\mu_{S^{\prime},1}-\frac{1}{\#\mathbf{O}(\mathbf{x})}\sum_{\mathbf{y}\in\mathbf{O}(\mathbf{x})}f_{S^{\prime},1}(\mathbf{y})\right|<\varepsilon_{S^{\prime}}$
for all $\mathbf{x}\in(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\setminus Z_{S^{\prime},1})(\overline{\mathbb{Q}})$ with $\hat{h}_{{\mathcal{L}}^{\boxtimes m(M+2)}}(\mathbf{x})\geq\delta_{\varepsilon_{S^{\prime}},1}$; here $\mathbf{O}(\mathbf{x})$ is the Galois orbit of $\mathbf{x}$ (over $\mathbb{Q}$). Applying the lemma to $\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})$, $f_{S^{\prime},2}$ and $\varepsilon_{S^{\prime}}$, we obtain similarly a constant $\delta_{\varepsilon_{S^{\prime}},2}>0$ and a Zariski closed subset $Z_{S^{\prime},2}\subsetneq\mathscr{D}(\mathscr{X}_{S^{\prime}}^{[m(M+2)]})$ such that
@@
(5.9) $\left|\int_{\mathscr{D}(X_{S^{\prime}}^{[m(M+2)]})(\mathbb{C})}f_{S^{\prime},2}\mu_{S^{\prime},2}-\frac{1}{\#\mathbf{O}(\mathbf{x}^{\prime})}\sum_{\mathbf{y}^{\prime}\in\mathbf{O}(\mathbf{x}^{\prime})}f_{S^{\prime},2}(\mathbf{y}^{\prime})\right|<\varepsilon_{S^{\prime}}$
for all $\mathbf{x}^{\prime}\in(\mathscr{D}(\mathscr{X}_{S^{\prime}}^{[m(M+2)]})\setminus Z_{S^{\prime},2})(\overline{\mathbb{Q}})$ with $\hat{h}_{{\mathcal{L}}^{\boxtimes m(M+1)}}(\mathbf{x}^{\prime})\geq\delta_{\varepsilon_{S^{\prime}},2}$.
Set $\delta_{S^{\prime}}:=\min\{\delta_{\varepsilon_{S^{\prime}},1},\delta_{\varepsilon_{S^{\prime}},2}\}>0$ and define the Zariski closed set $Z_{S^{\prime}}=Z_{S^{\prime},1}\cup\mathscr{D}^{-1}(Z_{S^{\prime},2})$. Note that $Z_{S^{\prime}}\neq\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$ as $\mathscr{D}$ is generically finite.
@@
Claim: For each point $\mathbf{x}\in(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\setminus Z_{S^{\prime}})(\overline{\mathbb{Q}})$, we have